There are 20 machines in a large plant. The machines operate independently of one another, breakdowns occur at random. The probability of a breakdown for each machine in one day’s operation is 0.10. Find the probability that exactly 4 breakdowns occur in one day


1
Expert's answer
2021-03-01T17:12:21-0500

By condition,


"p = 0.1 \Rightarrow q = 1 - p = 0.9"


Then, according to Bernoulli's formula, the wanted probability is


"{P_{20}}\left( 4 \right) = C_{20}^4{p^4}{q^{20 - 4}} = \frac{{20!}}{{4!16!}} \cdot {0.1^4} \cdot {0.9^{16}} =\\= \frac{{17 \cdot 18 \cdot 19 \cdot 20}}{{24}} \cdot {0.1^4} \cdot {0.9^{16}} \approx 0.0898"


Answer: "\approx 0.0898."